OpenStudy (republican31):

Solve 3 – 2(b – 5) = 7. Or is the equation an Identity? Or if it has any solution at all? I want help to solve it more than the straight answer.

OpenStudy (anonymous):

so you just want to solve for b right?

OpenStudy (republican31):

Yes. If it is solvable.

OpenStudy (anonymous):

okay so you can start of by distributing the -2 into (b - 5) do you know how to do that?

OpenStudy (rhr12):

b=12

OpenStudy (republican31):

Yes. 3-2b+10=7, right?

OpenStudy (anonymous):

yes okay so now you can combine the 3 and the 10 into a 13

OpenStudy (anonymous):

also b does not =12 @rhr12

OpenStudy (anonymous):

okay so now you should have -2b + 13 = 7

OpenStudy (republican31):

Okay so 13-2b=7. Then subtract 13 from both sides?

OpenStudy (rhr12):

(3-2)*{b(12)-5}=7 1*7=7 7=7

OpenStudy (rhr12):

i used a different formula but i also like your formula @watchpencilpaper

OpenStudy (republican31):

-2b = -6. So then divide -6 by -2, making b=3, correct?

OpenStudy (republican31):

@rhr12 that would make it an identity, correct? If it is correct, @watchpencilpaper?

OpenStudy (republican31):

If the equation is correct, I mean.

OpenStudy (anonymous):

you cant subtract the 2 from the 3 before multiplying it by the (12-5) so its wrong

OpenStudy (anonymous):

lets just continue solving the equation and then we can check our answer so we left of at -2b = -6 right?

OpenStudy (republican31):

Yes. b=3?

OpenStudy (anonymous):

yes okay and we can check that by plugging it back into our equation \[3 - 2(b-5) = 7\] \[3 - 2((3)-5) = 7\] \[3 - 2(-2) = 7\] \[3 - (-4) = 7\] \[3 + 4 = 7\] \[7 = 7\]

OpenStudy (republican31):

Alright! Thank you!

OpenStudy (anonymous):

yep no problem :D